# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v1

@@ Abstract
This work presents FOOL, an OEC-native and task-agnostic feature compression method that preserves prediction performance. FOOL partitions high-resolution satellite imagery to maximize throughput. Further, it embeds context and leverages inter-tile dependencies to lower transfer costs with negligible overhead. While FOOL is a feature compressor, it can recover images with competitive scores on perceptual quality measures at lower bitrates. We extensively evaluate transfer cost reduction by including the peculiarity of intermittently available network connections in low earth orbit. Lastly, we test the feasibility of our system for standardized nanosatellite form factors. We demonstrate that FOOL permits downlinking over 100$\times$ the data volume without relying on prior information on the downstream tasks.
@@ 1 Introduction
As constellation sizes and sensor resolutions increase, downlink bandwidth cannot keep up with the accumulating data volume [7, 8]. Although additional ground station equipment may prevent link saturation, building and maintaining them, including licensing the necessary frequencies, is a significant cost factor for satellite operation. As an alternative, Orbital Edge Computing (OEC) proposes processing data at the source [9, 10, 11, 12, 13]. Recent work on reducing bandwidth requirements in OEC is roughly categorizable in aggressive (task-oriented) filtering and compression [14]. The former relies on subjective value measures that restrict their practicability to coarse-grained tasks, such as de-duplication or cloud filtering. The latter constrains entire missions to particular tasks or prediction models. We argue that the limitations of existing compression or other data reduction approaches are particularly adverse to OEC.
@@
First, the CubeSat design is intended for short-duration missions [3] (typically up to 3-5 years), and despite waning prices, launching sensor networks in space is still associated with substantial logistical, administrative, and monetary costs. Therefore, it seems undesirable to designate entire constellations to a small subset of tasks and Deep Neural Network (DNN) architectures. More pressingly, irrespective of whether current codecs can prevent bottlenecks, they may undermine the effectiveness of entire missions. Precisely, the assumption of prediction models (i.e., machine perception) only requiring a subset of information for image reconstruction (i.e., human perception) may lead to false confidence in a codec to reliably discern the salient signals. We argue the opposite holds, i.e., when the objective is to accommodate arbitrary downstream tasks with prediction models instead of human experts, there is less potential for rate reductions. Intuitively, two seemingly visually identical images may have subtle differences in pixel intensities, which a prediction model could leverage to overcome physiological restrictions.
@@
In summary, three conflicting objectives aggravate the challenges for OEC: (i) maximizing downlinking captures, (ii) ensuring the value of the captures by relying on as few assumptions on downstream tasks as possible, and (iii) minimizing the risk from unpredictable adverse effects on current and future prediction models. To this end, we propose drawing from recent work on neural feature compression with Shallow Variational Bottleneck Injection (SVBI) [15, 16, 17]. The idea of SVBI is to reduce the risk of discarding information necessary for arbitrary, practically relevant tasks by targeting the shallow representation of foundational models as a reconstruction target in the rate-distortion objective. In other words, rate gains come from constraining the solution space with abstract high-level criteria rather than reifying target tasks with an explicit definition of value or expert-crafted labels. We investigate whether the SVBI framework is suitable for EO from a compression perspective and identify lower-level system considerations given the oppressive constraints of OEC. Then, we apply our insights to introduce a Tile Holistic Efficient Featured Oriented Orbital Learned (THE FOOL) compression method, which we will refer to as FOOL for short. FOOL alleviates the challenges of OEC by generalizing SVBI to improve compression performance while introducing more specific methods that aid in meeting the requirements of OEC and EO tasks. FOOL comprises a simple pipeline, a profiler, and a neural feature codec with a separate reconstruction model. The pipeline minimizes overhead by CPU-bound pre- and post-processing with concurrent task execution. The profiler identifies configurations that maximize data size reduction, factoring in intermittently available downlinks and the trade-off between processing throughput and lowering bitrate from more powerful but costlier transforms. The neural codec’s architecture includes task-agnostic context and synergizes with the profiler’s objective to maximize throughput with batch parallelization by exploiting inter-tile spatial dependencies.
@@ 2.2 Preventing Link Saturation with Orbital Inference
Gadre et al. introduce Vista [24], a Joint Source Channel Coding (JSCC) system for LoRa-enabled CubeSats, designed to enhance low-latency downlink communication of satellite imagery and DNN inference. It shows significant improvements in image quality and classification performance through LoRa-channel-aware image encoding. Moreover, the evaluation assumes simple tasks that are not representative of practical EO. In contrast, FOOL decouples image recovery from the initial compression objective and ensures task-agnostic preservation of information.
Lu et al. introduce STCOD [25], a JSCC system for efficient data transmission and object detection in optical remote sensing. STCOD integrates satellite computing to process images in space, distinguishing between regions of interest (ROIs) and backgrounds. It shows promising results with a block-based adaptive sampling method, prioritizing transmitting valuable image blocks using fountain code [26]. The caveat is that ROI detectors that can reliably prevent predictive loss for downstream tasks require strong biases regarding sensor properties and task granularity. FOOL includes task-agnostic context, with significantly less overhead and more robustness towards varying conditions than an ROI detector. Furthermore, it is end-to-end optimized with the other compression model components without relying on the same biases or expected-crafted labels.
Thematically, our work resembles Kodan by Denby et al. [27] the closest. Like FOOL, Kodan treats channel conditions as an orthogonal problem and primarily focuses on source coding to address the downlink and computational bottlenecks. Kodan uses a reference application for satellite data analysis and a representative dataset to create specialized small models. Once in orbit, it dynamically selects the best models for each data sample to maximize the value of data transmitted within computational limitations. Kodan’s excellent system design is promising but relies on assumptions that hinder practicability and the potential for meaningful rate reductions. Conversely, FOOL follows a different design philosophy by treating the downlink bottleneck primarily as a compression problem. Like dimensionality reduction [28], filtering is only a crude approximation to compression, not equivalent to it. Additionally, we do not treat the computational deadline as a hard temporal constraint to decouple the method to a particular system design. Instead, FOOL’s profiler measures key performance indicators on the pixel level. Specifically, given hardware limitations, the aim is to reduce transfer cost by balancing the lower bitrate of more powerful encoders and the gain in processing throughput of more lightweight encoders.
@@ 3 Background & Problem Formulation
3.1 Computation at the Space’s Edge
Unlike terrestrial edge computing, OEC is subject to additional oppressive limitations on resources for processing and communication. Besides volatile link conditions, methods must consider the constraints of LEO (nano-)satellites, usually operating at altitudes from 300 to 800km to reduce launch costs and increase imaging resolution [9]. A 3U nanosatellite following the CubeSat standard is limited to 10cm×$\times$10cm×$\times$30cm and 4kg [29] with restricted power supply by using solar harvesting [30].
@@ 3.2 Limitations of Codecs
Given remote image captures and a set of unknown associated object detection tasks, we seek a transformation of the sensings into representations that minimize transfer costs and loss of information that may impact any detection tasks (i.e., predictive loss) without relying on strong assumptions. For the remainder of this work, we refer to generalizability as measuring how well a method can minimize the predictive loss on unknown (i.e., without prior information) detection tasks. For example, assume a client is interested in detecting large vehicles $Y_{vh}$, and their model achieves $0.65$ mean average precision (mAP). We could choose the identity function for our transformation, generalizing to all tasks, i.e., trivially resulting in no predictive loss for $Y_{vh}$. In contrast, task-oriented transforms (e.g., [31]) retain information for $Y_{vh}$. Although this may achieve lossless prediction for $Y_{vh}$, it does not generalize, as the transformed data is unusable for non-overlapping tasks, presenting a considerable disadvantage for FAIR [32] data reuse.
3.2.1 Handcrafted and Learned Image Codecs
An intuitive approach is to apply lossless codecs. Besides bent pipes, it is the only approach with easily understood guarantees on generalization. Nevertheless, lossless compression cannot adequately address the downlink bottleneck due to theoretical lower bounds. More promising alternatives are lossy methods that require relaxing the requirement of relying on identical reconstruction for generalization. More formally, given a distortion measure $\mathcal{D}$, a constraint $D_{c}$ bounds the minimal bitrate to [33]:
@@ 3.2.1 Handcrafted and Learned Image Codecs
We download pre-trained weights [36] for the ImageNet [37] classification task of three popular architectures[38, 39, 40]. First, we compute the expected Peak Signal-To-Noise Ratio (PSNR) of a popular LIC model [41] for each quality level on the validation set. Then, we apply Additive White Gaussian Noise (AWGN) on input to match the PSNR of a codec for each quality level separately. Lastly, we measure the predictive loss as the average difference between the accuracy of the original and processed samples.
Notice that the predictive loss on the distorted input is significantly worse than the noisy input. Additive noise does not remove information; rather, it superimposes unwanted information. Conversely, lossy compression intentionally discards information from signals, and two codecs may achieve comparable rate-distortion performance despite emphasizing different information to retain. Re-training model weights on reconstructed samples may mitigate some predictive loss, but only due to adjusting prediction to input perturbations (e.g., quantization artifacts) and error-prone extrapolation of lost information.
Particularly, for EO with satellite imagery that spans large geographic areas we stress the unsuspecting danger of lossy compression which is compounded with learned transforms [21], where it is challenging to understand behavior. The ability to differentiate between intensities beyond the capability of human perception may explain why detection models can outperform human experts. Accordingly, we should assume that lossy codecs may discard information where even experts cannot reliably verify the impact on machine interpretability. For example, suppose a codec that reduces the rate by focusing on preserving coarser-grained structures. Then, tasks that rely on assessing the environment for fine-grained object classes will lack background information (e.g., inferring region by tree species with subtle color variations) for correct predictions.
@@
Current limitations of image codecs put operators in a difficult position, especially for EO. The decision falls between (a combination of) lossless codecs, applying crude filtering methods, or attempting to reduce the bitrate with lossy codecs, remaining uncertain about whether the codecs retain information necessary for detection models in real-life conditions. As a solution, we advocate for Shallow Variational Bottleneck Injection (SVBI) [17], which uses shallow features as the reconstruction target that prioritizes salient regions for (near) arbitrary high-level vision tasks.
@@ 3.3 Shallow Variational Bottleneck Injection
SVBI trains neural codecs by replacing the distortion term of the rate-distortion objective of variational image compression models [34] with head distillation (HD) [42, 43, 15, 16, 17]. Figure 2 illustrates an example, where the HD distortion measure penalizes a compression model for not sufficiently approximating the shallow representation of a pre-trained feature extractor.
In Knowledge Distillation (KD) terminology, the codec is referred to as the student and the shallow layers of a foundational model as the teacher. Note that KD is not this work’s focus, as HD diverges from the typical KD objective. The intuition behind SVBI is that if a codec can reconstruct the representation of a foundational model, then the representation is sufficient for at least all tasks associated with that model.

### v2

@@ Abstract
This work presents FOOL, an OEC-native and task-agnostic feature compression method that preserves prediction performance. FOOL partitions high-resolution satellite imagery to maximize throughput. Further, it embeds context and leverages inter-tile dependencies to lower transfer costs with negligible overhead. While FOOL is a feature compressor, it can recover images with competitive scores on quality measures at lower bitrates. We extensively evaluate transfer cost reduction by including the peculiarity of intermittently available network connections in low earth orbit. Lastly, we test the feasibility of our system for standardized nanosatellite form factors. We demonstrate that FOOL permits downlinking over 100$\times$ the data volume without relying on prior information on the downstream tasks.
@@ 1 Introduction
As constellation sizes and sensor resolutions increase, downlink bandwidth cannot keep up with the accumulating data volume [7, 8]. Additional ground station equipment may prevent link saturation, building and maintaining them, including licensing the necessary frequencies, is a significant cost factor for satellite operation. As an alternative, Orbital Edge Computing (OEC) proposes processing data at the source [9, 10, 11, 12, 13]. Recent work on reducing bandwidth requirements in OEC is roughly categorizable in aggressive (task-oriented) filtering and compression [14]. The former relies on subjective value measures that restrict their practicability to coarse-grained tasks, such as de-duplication or cloud filtering. The latter constrains entire missions to particular tasks or prediction models. We argue that the limitations of existing compression or other data reduction approaches are particularly adverse to OEC.
@@
First, the CubeSat design is intended for short-duration missions [3] (typically up to 3-5 years), and despite waning prices, launching sensor networks in space is still associated with substantial logistical, administrative, and monetary costs. Therefore, it seems undesirable to designate entire constellations to a small subset of tasks and Deep Neural Network (DNN) architectures. More pressingly, irrespective of whether current codecs can prevent bottlenecks, they may undermine the effectiveness of entire missions. Precisely, the assumption of prediction models only requiring a subset of information for image reconstruction may lead to false confidence in a codec to reliably discern the salient signals. We argue the opposite holds, i.e., when the objective is to accommodate arbitrary downstream tasks with prediction models instead of human experts, there is less potential for rate reductions. Intuitively, two seemingly visually identical images may have subtle differences in pixel intensities, which a prediction model could leverage to overcome physiological restrictions.
@@
In summary, three conflicting objectives aggravate the challenges for OEC: (i) maximizing downlinking captures, (ii) ensuring the value of the captures by relying on as few assumptions on downstream tasks as possible, and (iii) minimizing the risk from unpredictable adverse effects on current and future prediction models. To this end, we propose drawing from recent work on neural feature compression with Shallow Variational Bottleneck Injection (SVBI) [15, 16, 17]. The idea of SVBI is to reduce discarding information necessary for arbitrary, practically relevant tasks by targeting the shallow representation of foundational models as a reconstruction target in the rate-distortion objective. In other words, rate reductions come from constraining the solution space with abstract high-level criteria rather than reifying target tasks with an explicit definition of value or expert-crafted labels. We investigate whether the SVBI framework is suitable for EO from a compression perspective and identify lower-level system considerations given the oppressive constraints of OEC. Then, we apply our insights to introduce a Tile Holistic Efficient Featured Oriented Orbital Learned (THE FOOL) compression method, which we will refer to as FOOL for short. FOOL alleviates the challenges of OEC by generalizing SVBI to improve compression performance while introducing more specific methods that aid in meeting the requirements of OEC and EO tasks. FOOL comprises a simple pipeline, a profiler, and a neural feature codec with a separate reconstruction model. The pipeline minimizes overhead by CPU-bound pre- and post-processing with concurrent task execution. The profiler identifies configurations that maximize data size reduction, factoring in intermittently available downlinks and the trade-off between processing throughput and lowering bitrate from more powerful but costlier transforms. The neural codec’s architecture includes task-agnostic context and synergizes with the profiler’s objective to maximize throughput with batch parallelization by exploiting inter-tile spatial dependencies.
@@ 2.2 Preventing Link Saturation with Orbital Inference
Gadre et al. introduce Vista [24], a Joint Source Channel Coding (JSCC) system for LoRa-enabled CubeSats, designed to enhance low-latency downlink communication of satellite imagery and DNN inference. It shows significant improvements in image quality and classification performance through LoRa-channel-aware image encoding. Moreover, the evaluation assumes simple tasks not representative of practical EO. In contrast, FOOL decouples image recovery from the initial compression objective and ensures task-agnostic preservation of information.
Lu et al. introduce STCOD [25], a JSCC system for efficient data transmission and object detection in optical remote sensing. STCOD integrates satellite computing to process images in space, distinguishing between regions of interest (ROIs) and backgrounds. It shows promising results with a block-based adaptive sampling method, prioritizing transmitting valuable image blocks using fountain code [26]. The caveat is that ROI detectors that can reliably prevent predictive loss for downstream tasks require strong biases regarding sensor properties and task granularity. FOOL includes task-agnostic context, with significantly less overhead and more robustness towards varying conditions than an ROI detector. Furthermore, it is end-to-end optimized with the other compression model components without relying on the same biases or expert-crafted labels.
Thematically, our work resembles Kodan by Denby et al. [27] the closest. Like FOOL, Kodan treats channel conditions as an orthogonal problem and primarily focuses on source coding to address the downlink and computational bottlenecks. Kodan uses a reference application for satellite data analysis and a representative dataset to create specialized small models. Once in orbit, it dynamically selects the best models for each data sample to maximize the value of data transmitted within computational limitations. Kodan’s excellent system design is promising but relies on assumptions that hinder practicability and the potential for meaningful rate reductions. Conversely, FOOL follows a different design philosophy by treating the downlink bottleneck primarily as a compression problem. We do not treat the computational deadline as a hard temporal constraint to decouple the method to a particular system design. FOOL’s profiler measures key performance indicators on the pixel level. Given hardware limitations, the aim is to reduce transfer cost by balancing the lower bitrate of more powerful encoders and the gain in processing throughput of more lightweight encoders.
@@ 3 Background & Problem Formulation
3.1 The Downlink Bottleneck
Downlink bottlenecks occur when the data volume exceeds the bandwidth within a downlink segment during a single pass. We formalize a model sufficient for our purposes by considering link conditions and sensor properties of satellites belonging to a constellation. A constellation is defined as $\mathcal{C}=(L,\mathcal{S},I,f)$ where $L$ is a link to communicate with a ground station and $\mathcal{S}$ is a set of satellites. The link is determined by its expected downlink rate, measured in megabits per second (Mbps). The function $f:S\rightarrow I$ maps each satellite $s\in S$ to an interval where it passes the downlink segment into disjoint subsets $\mathcal{G}=\{\mathcal{G}_{i}|\mathcal{G}_{i}=\{s\in\mathcal{S}|f(s)=i\},i\in I\}$, such that $\bigcup\mathcal{G}_{i}=\mathcal{S}$ and $\bigcap\mathcal{G}_{i}=\varnothing$. The link capacity $V_{\text{link}}$ is the bandwidth available per pass and is determined by the link rate and the interval range. Satellites $S=(R_{\text{orbit}},S_{\text{rate}},S_{\text{spatial}},S_{\text{bands}},S_{\text{radio}},S_{\text{fov}})$ are equipped with a sensor, and its properties determine the volume per capture.
$V_{\text{capture}}=\overbrace{\frac{R^{2}_{\text{orbit}}\cdot\operatorname{tan}^{2}(S_{\text{fov}})}{S_{w}\cdot S_{h}}}^{\text{Total Pixels}}\cdot\underbrace{S_{\text{bands}}\cdot S_{\text{radio}}}_{\text{Bits per Pixel}}$ (1)
The radiometric resolution $S_{\text{radio}}$ and number of bands $S_{\text{bands}}$ determine the downlink cost per pixel in bits. The orbit $R_{\text{orbit}}$, sensor spatial resolution $S_{\text{spatial}}=S_{h}\times S_{w}$, and field of view $S_{\text{fov}}$ determine the number of pixels per capture. The number of captures depends on the time to complete an orbit
$T_{\text{orbit}}=2\pi\sqrt{\frac{(R_{\text{orbit}}+R_{\text{earth}})^{3}}{GM}}$ (2)
and on the capture rate $S_{\text{rate}}$. $G$ is the gravitational constant and $M$ is the earth’s mass. The orbit $R_{\text{orbit}}$ is usually around 160 to 800 kilometers for LEO satellites. For reference, $R_{\text{orbit}}$ is 786 kilometers for Sentinel-2 [28]. Finally, the number of captures from all satellites within the segmentation group determines the total volume per pass.
$V_{\text{pass}}=\sum_{s^{(i)}\in\mathcal{G}_{j}}T_{\text{Orbit}}\cdot S^{(i)}_{\text{rate}}\cdot V^{(i)}_{\text{capture}}$ (3)
The superscript $(i)$ denotes the costs associated with a satellite $(i)$. For constellations with homogenous sensors $V_{\text{capture}}$ is a static value. Notice that $V_{\text{pass}}$ scales linearly by the constellation size and a constant factor $c$ for overlap occurrences, i.e., $|\mathcal{G}_{j}|=\frac{|\mathcal{S}|}{c}$. To determine $c$ for a constellation, we must calculate the minimum angle between satellites $\beta^{*}$. Assuming a single ground station at the Earth’s North Pole and given the minimum communication elevation $\theta$
$\beta^{*}=2\times(180^{\circ}-(\theta+90^{\circ})-\arcsin({\frac{R_{\text{earth}}\cdot\sin{(90^{\circ}+\theta)}}{R_{\text{orbit}}+R_{\text{earth}}}})$ (4)
For example, consider a constellation at $R_{\text{orbit}}=790,000$ meters altitude with a minimum elevation $\theta=25^{\circ}$ such that $\beta^{*}\approx 22.52^{\circ}$. Then, $c=\frac{360^{\circ}}{22.52^{\circ}}\approx 16$, i.e., to prevent any interval sharing, the constellation size may not exceed 16 satellites.
In short, the aim is to facilitate cost-efficient scaling of constellations by increasing bandwidth value and substantially reducing reliance on building additional infrastructure. That is, we require an encoding scheme enc, such that $V_{\texttt{enc}}<V_{\text{link}}$. Note that a single satellite may experience a bottleneck even if the constellation is sparse enough to prevent interval sharing [27]. Say, each $s\in\mathcal{S}$ is equipped with a sensor using approximate Sentinel-2 configurations [28] by setting a multispectral sensor for (near-) visible light to $S_{\text{bands}}=4$, $S_{\text{radio}}=12$ $S_{\text{fov}}=21^{\circ}$, $S_{h\times w}=10\times 10$, and five captures per pass. With $|\mathcal{S}|\leq 16$, the volume for each pass is $\frac{790,000^{2}\cdot\operatorname{tan}^{2}(10.5^{\circ})}{100}\cdot 4\cdot 12\cdot 5\approx 410$ GB. To prevent a bottleneck even without sharing an interval and using a higher-end link, such as WorldView-3 [29] where $V_{\text{link}}=90$ GB per pass, the enc needs to decrease the data volume by a factor of 4.5.
There are two overarching objectives for a codec and the system in which we deploy its encoder. The system’s objective is to process and encode large volumes of high-dimensional data, given the physical limitations of LEO (nano-) satellites. A 3U nanosatellite following the CubeSat standard is limited to 10cm×$\times$10cm×$\times$30cm and 4kg [30] with restricted power supply by using solar harvesting [31]. The compression objective is to achieve a sufficiently low bitrate while maintaining the data’s integrity. The following elaborates on the challenges of conceiving a method that fulfills our criteria and the limitations of applying existing codecs.
@@ 3.2 Limitations of Codecs
Given remote image captures and a set of unknown associated object detection tasks, we seek a transformation of the captures into representations that minimize transfer costs and loss of information that may impact any detection tasks. We refer to generalizability as a measure of how well a method can minimize the predictive loss on unknown detection tasks. For example, a purely task-oriented encoding (e.g., [32]) can retain information for a set of explicitly defined tasks. Still, it does not generalize as the transformed data is unusable for non-overlapping tasks Besides bent pipes, textitlossless codecs are the only approach with easily understood guarantees on generalization. Nevertheless, lossless compression cannot adequately address the downlink bottleneck due to theoretical lower bounds. Promising alternatives are lossy methods that relax the requirement of relying on identical reconstruction for generalization. More formally, given a distortion measure $\mathcal{D}$, a constraint $D_{c}$ bounds the minimal bitrate to [33]:
@@
We download pre-trained weights [37] for the ImageNet [38] classification task of three popular architectures[39, 40, 41]. First, we compute the expected Peak Signal-To-Noise Ratio (PSNR) of a popular LIC model [42] for each quality level on the validation set. Then, we apply Additive White Gaussian Noise (AWGN) on input to match the PSNR of a codec for each quality level separately. Lastly, we measure the predictive loss as the average difference between the accuracy of the original and processed samples. Notice that the predictive loss on the distorted input is significantly worse than the noisy input. Additive noise does not remove information; rather, it superimposes unwanted information. Conversely, lossy compression intentionally discards information from signals, and two codecs may achieve comparable rate-distortion performance despite emphasizing different information to retain. Re-training model weights on reconstructed samples may mitigate some predictive loss, but only due to adjusting prediction to input perturbations and error-prone extrapolation of lost information.
Particularly, for EO with satellite imagery that spans large geographic areas, we stress the unsuspecting danger of lossy compression, which is compounded with learned transforms [21], where it is challenging to understand behavior. The ability to differentiate between intensities beyond the capability of humans may explain why detection models can outperform domain experts. Accordingly, we should assume that lossy codecs may discard information where even experts cannot reliably verify the impact on machine interpretability. For example, suppose a codec that reduces the rate by focusing on preserving coarser-grained structures. Then, tasks that rely on assessing the environment for fine-grained object classes will lack background information (e.g., inferring region by tree species with subtle color variations).
@@
Current limitations of image codecs put operators in a difficult position, especially for EO. The decision falls between (a combination of) lossless codecs, applying crude filtering methods, or attempting to reduce the bitrate with lossy codecs, remaining uncertain about whether the codecs retain information necessary in real conditions. As a solution, we advocate for Shallow Variational Bottleneck Injection (SVBI) [17], which prioritizes salient regions for (near) arbitrary high-level vision tasks.
@@ 3.3 Shallow Variational Bottleneck Injection
SVBI trains neural codecs by replacing the distortion term of the rate-distortion objective of variational image compression models [34] with head distillation (HD) [43, 44, 15, 16, 17]. Figure 2 illustrates an example, where the HD distortion measure penalizes a compression model for not sufficiently approximating the shallow representation of a pre-trained foundational model. We define a foundational model as a pre-trained DNN that can accommodate multiple tasks by attaching predictors or fine-tuning the deeper layers. In Knowledge Distillation (KD) terminology, the codec is referred to as the student and the shallow layers of a foundational model as the teacher. Note that KD is not this work’s focus, as HD diverges from the typical KD objective. The intuition behind SVBI is that if a codec can reconstruct the representation of a foundational model, then the representation is sufficient for at least all tasks associated with that model.